Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 30/8/2017

Location : CR21 at 14:00 pm

Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef): completed draft exists but further changes expected; stands
    • Update shift checklist to focus on alarms : (Steve): stands (need to sit down with Paolo)
    • Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands
  • DC update
  • Hydrogen system update and plans (JBoehm)
  • Status of alarm clearance (DC/Ajit)
    • Update to search procedure documentation for hydrogen system (Craig/Henry) - in progress
  • Update on tracker readout (Melissa)
  • Update on SS QPS connector issue (Craig)
  • Cycle 2017/01 Issues
  • Hall/Operations plan : 17-08-17-CW.xlsx
  • Operations startup
    • Initial discussion on system readiness : Wednesday August 23rd in CR25 in R1 @ 11
    • Pre-startup meeting : Monday Sep 4th (Postponed until after the pulser run)
    • Pulser run : Sep 7th/8th
    • Outcome of ISIS machine physics scheduling meeting
      • MICE has tentatavely been allocated 08:30 - 16:30 on Sept 14th
      • ISIS would not be surprised if they lost a dipole during the user run-
        sounds like something to do with air between the coils and the pipework. If things don't go
        smoothly then likely MICE will be rescheduled.
  • Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
    No.
  • AOB

Minutes

  • Actions
    • Give LH2 shifter documentation to Steve to add to official documentation list (Josef): completed draft exists but further changes expected from Josef & Andy; stands
    • Update to search procedure documentation for hydrogen system (Craig/Henry) - Done.
    • Update shift checklist to focus on alarms : (Steve): stands Aims to make progress now Paolo is back
    • Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands
  • DC update

    Target Frame test ok, Dipoles ok, Diffuser ok & Proton Ab ok. Quads test imminent.
    A/C unit 4 now also has an alarm, about changing the filter. Running at risk. Appreciate we need some maintenance. CW try and catch Paul Masterson.
    Also Paul M is needed to ensure SS fall-over cooling is in place

  • Status of alarm clearance (DC/Ajit) - No update
  • Hydrogen system update and plans (JBoehm)
    Vent line leak checking in progress
    Paolo setting up Alarm handler with beep in ISIS LCR
    Need H2 bottle pack in cage - AN overseeing this coming Thursday morn.
    If we get safety tour sign-off then we will cool with He the rest of this week and then system proper next week with H2
    Tour starts 10am tomorrow in MLCR
  • Operations startup
    • Pulser run : Sept 7th. Melissa, Agit and Durga to attend.
    • Beam tuning and beam test run : sometime in the week starting Sep 11th. Will ask ISIS for 8 hours during this week.
      CRogers reported that ISIS will run at risk at 800MeV and could be susceptible to dipole problem. If they have to reduce to 700MeV then it will take a few days to sort out.
  • Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
    No.
  • AOB
    DS is ok but unstable. Shield supply pressure is wavering. Mark T to investigate when freed from LH2 responsibilities.

Updated by Boyd, Steven about 5 years ago · 10 revisions